# Functional Graphical Models: Structure Enables Offline Data-Driven Optimization

### v2

@@ Appendix C Details of Experiments
In all runs, all neural networks were two-hidden layer MLPs, with 256 hidden units. Each of them was trained with Adam (Kingma & Ba, 2014) optimization algorithm with batch size of 128. Hyperparameters for COMs were taken from the COMs paper (Trabucco et al., 2021) and we did not observe any improvement by tuning them. As hyperparameters for all other methods are a subset of hyperparameters of COMs, we used the same hyperparameters, except that methods that use VAEs (Kingma & Welling, 2013) have 10 times smaller gradient ascent learning rate. We also used the common choice for temperature $\tau=0.05$ for reward-weighted regression (Peters & Schaal, 2007; Peng et al., 2019; Kostrikov et al., 2021).
@@
$\begin{bmatrix}{\bm{C}}_{11}{\mathbf{x}}_{1}\quad\dots\quad{\bm{C}}_{1d}{\mathbf{x}}_{d}\\ \\ \dots\\ \\ {\bm{C}}_{|\mathcal{C}|1}{\mathbf{x}}_{1}\quad\dots\quad{\bm{C}}_{|\mathcal{C}|d}{\mathbf{x}}_{d}\end{bmatrix}$ (21)

### v3

@@ Appendix C Details of Experiments
In all runs, all neural networks were two-hidden layer MLPs, with 256 hidden units. Each of them was trained with Adam (Kingma and Ba, 2014) optimization algorithm with batch size of 128. Hyperparameters for COMs were taken from the COMs paper (Trabucco et al., 2021) and we did not observe any improvement by tuning them. As hyperparameters for all other methods are a subset of hyperparameters of COMs, we used the same hyperparameters, except that methods that use VAEs (Kingma and Welling, 2013) have 10 times smaller gradient ascent learning rate. We also used the common choice for temperature $\tau=0.05$ for reward-weighted regression (Peters and Schaal, 2007; Peng et al., 2019; Kostrikov et al., 2021).
@@
$\begin{bmatrix}{\bm{C}}_{11}{\mathbf{x}}_{1}\quad\dots\quad{\bm{C}}_{1d}{\mathbf{x}}_{d}\\ \\ \dots\\ \\ {\bm{C}}_{|\mathcal{C}|1}{\mathbf{x}}_{1}\quad\dots\quad{\bm{C}}_{|\mathcal{C}|d}{\mathbf{x}}_{d}\end{bmatrix}$